OpenStudy (anonymous):

Mark is 3 times as old as Sheri. In 5 years, the difference in their ages will be 24 years. How old are they now?

OpenStudy (anonymous):

So we need to write out a couple of equations to solve this. Let M=Mark's age, and S=Sheri's age. M=3(S) (M+5) - (S+5) = 24 Do you understand how I got those equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but I am still confuse how to get the answer

OpenStudy (misty1212):

put the \(3S\) for \(M\) and solve \[3S+5-(S+5)=24\]

OpenStudy (anonymous):

That's fine, that was just the first step. Now you have two equations, M=3(S) (M+5) - (S+5) = 24 We can solve them. You can substitute M = 3(S) into the other equation: (3S + 5) - (S + 5) = 24 Now you just have to simplify that equation to find S

OpenStudy (anonymous):

s=12

OpenStudy (anonymous):

s=12 @livias.random

OpenStudy (anonymous):

Exactly. Now that you know what S is, substitute S = 12 into M = 3(S) and solve for M

OpenStudy (anonymous):

M=36

OpenStudy (anonymous):

Right, so now you have both ages :)